OpenStudy (anonymous):

Brandon owns a farm stand and mixes apple juice and cranberry juice. how much should he charge if he mixed 8L of apple juice selling for $0.45/L with 10L of cranberry juice for $1.08/L? can you help me out

OpenStudy (anonymous):

\[mix = \frac{ (8L \times $0.45/L + 10L \times $1.08 / L) }{18 L }$/L\] \[mix = \frac{ ($3.6 + $10.80) }{18 L }$/L\]

OpenStudy (anonymous):

do you need what he charge for the mix in $x.xx/L or what he charges in total for the 18L of mix?

OpenStudy (anonymous):

what he charges total

OpenStudy (anonymous):

in total just the numerator of the formula above: $3.6 for 8L x $0.45 plus $10.80 for 10L x $1.08

OpenStudy (anonymous):

thanks mathessentials

OpenStudy (anonymous):

you're welcome ;-)